OpenStudy (anonymous):

If (4xy^3 + 8x^2y^5)/( 2xy^2) is completely simplified to 2xayb + 4xcyd, where a, b, c, and d represent integer exponents, what is the value of a? Will fan and give medals

OpenStudy (anonymous):

@hareen

OpenStudy (anonymous):

hey dude

OpenStudy (anonymous):

what is up

OpenStudy (anonymous):

hey, can you help me?

OpenStudy (anonymous):

let me see

OpenStudy (anonymous):

its 7.10

OpenStudy (anonymous):

is this the pratice test

OpenStudy (anonymous):

nah its the dividing polynomals

OpenStudy (anonymous):

1 min brb

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

ok so what is answer that you got

OpenStudy (anonymous):

do you remeber the properties of expoents

OpenStudy (anonymous):

i got 3, i may be wrong, and yes

OpenStudy (anonymous):

let me check the answer

OpenStudy (anonymous):

when you like factor it completely you were supposed to get 2y • (2xy2 + 1)

OpenStudy (anonymous):

did you get that

OpenStudy (anonymous):

no, i just solved the problem, oh gosh im super wrong

OpenStudy (anonymous):

i meant divide

OpenStudy (anonymous):

so what did you get

OpenStudy (anonymous):

4xy^3 + 2y

OpenStudy (anonymous):

then what is a

OpenStudy (anonymous):

0?

OpenStudy (anonymous):

are you doing a quiz

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

are you in the new version

OpenStudy (anonymous):

like me

OpenStudy (anonymous):

whats the new version?

OpenStudy (anonymous):

it is like the for esol

OpenStudy (anonymous):

um i dont think so

OpenStudy (anonymous):

i meant like i took way long on segment 1 so it is like easy ersion

OpenStudy (anonymous):

no, im doing the reg one

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

yea i just not understanding this question

OpenStudy (anonymous):

ok i gotta go see ya later cause i need to like right now

OpenStudy (anonymous):

i know someone who can help

OpenStudy (anonymous):

@Mertsj

OpenStudy (anonymous):

okay see ya

OpenStudy (anonymous):

bye

OpenStudy (anonymous):

@Mertsj can you help?

OpenStudy (mertsj):

\[\frac{4xy^3+8x^2y^5}{2xy^2}=\frac{4xy^2(y+2xy^3)}{2xy^2}=2(y+2xy^3)=2y+4xy^3\]

OpenStudy (mertsj):

a=0

OpenStudy (anonymous):

thats what i thought, thank you

OpenStudy (mertsj):

yw